OpenStudy (anonymous):

How to find out endpoints? the midpoint of UV is (5,-11). the coordinates of one endpoint are U(3,5). find the coordinates of endpoint V. How do i figure this out when i know the midpoint?:(

OpenStudy (anonymous):

Use the distance formula to find the distance from the midpoint to the endpoint that you know, then add that distance to the midpoint in the opposite direction.

OpenStudy (anonymous):

ok:) what is the distance formula?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x_m = \frac{ x_1 + x_2 }{ 2 } \text{, } y_m = \frac{ y_1 + y_2 }{ 2 }\] you know the midpoint and an endpoint. put those in and solve for one you want

OpenStudy (anonymous):

don't do distance formula... long, tedious and unnecessary